OpenStudy (wheelchairgirl):

I really need help solving this and can you show all steps... (2√5+3√7)^2

OpenStudy (mertsj):

\[(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})\]

OpenStudy (mertsj):

Use FOIL

OpenStudy (blurbendy):

2sqrt[5] * 2sqrt[5] = 4sqrt[25] = 4 * 5 = 20 2sqrt[5] * 3sqrt[7] = 6sqrt[35] 2sqrt[5] * 3sqrt[7] = 6sqrt[35] 3sqrt[7] * 3sqrt[7] = 9sqrt[49] = 9 * 7 = 63 20 + 6sqrt[35] + 6sqrt[35] + 63 = 83 + 12sqrt[35]